OpenStudy (anonymous):

At noon, ship a is 150km west of ship b. ship a is sailing east at 35 km/h and ship b is sailing north at 25 km/h. how fast is the distance between the ships changing at 4 pm

OpenStudy (anonymous):

would i just take \[\Delta y / \Delta x \]

OpenStudy (amistre64):

you want dr/dt; so implicit this thing with respect to time.. your formula is the pythag thrm

OpenStudy (amistre64):

your dy/dt, or rather the d'north'/dt = 25; and d'e'/dt = 35

OpenStudy (amistre64):

e^2 + n^2 = r^2

OpenStudy (amistre64):

e' 2e + n' 2n = r' 2r ; divide it all by 2 e' e + n' n = r' r r' = [e' e + n' n]/ r

OpenStudy (anonymous):

ok so if i take that and introduce log i have f(x)+f'(x) which is

OpenStudy (amistre64):

log?

OpenStudy (anonymous):

\[e^2+n^2-r^2+ 1/2x +1/2n +1/2r\]

OpenStudy (anonymous):

well ln

OpenStudy (amistre64):

draw a triangle a right triangle with sides: n = 4(25); e = 4(35)+150; solve the pythag thrm for r r = sqrt(100^2 + e^2) :)

OpenStudy (amistre64):

there is no call for any log or ln in this problem... your thinking to much at it :)

OpenStudy (anonymous):

ok i am agreeing with using the pythag but i question your 4(35)+15 because of the direction of sail for the boat

OpenStudy (anonymous):

well 4(35)+150

OpenStudy (amistre64):

\[r' = \frac{35(e) + 25(n)}{\sqrt{100^2 + e^2}}\]

OpenStudy (anonymous):

because the boat is to the west and sailing east wouldn't it be 4(35)-150?

OpenStudy (anonymous):

ok let me work that out really fast

OpenStudy (amistre64):

150 + 4hours worth of 35 is.....150 + 4(35) :)

OpenStudy (amistre64):

....maybe, i coulda read it wrong lol

OpenStudy (anonymous):

ok for your r' i am showing 140.9996308

OpenStudy (amistre64):

your right; 4(35) - 150 :)

OpenStudy (amistre64):

2850 ------ sqrt(100^2 + 10^2) = sqrt(10100) r' = 28.358 is what I get

OpenStudy (anonymous):

ok if i use the reg pythag a^2+b^2=c^2 will i end up with the right answer after i solve the correct values

OpenStudy (amistre64):

i just renamed the variables to keep track of them better; I know east is 35 and n is 25, so to keep the variables in shape I just rename them n and e

OpenStudy (anonymous):

ok i did the same thing as you and ended up with a different number i am trying to work back to your answer

OpenStudy (amistre64):

n = 100; n' = 25; n n' = 2500 e = 10; e' = 25; e e' = 250

OpenStudy (amistre64):

r = sqrt(100^2 + 10^2) r = sqrt(10000 + 100) r = sqrt(10100)

OpenStudy (anonymous):

when i do sqrt(10100) i end up with 100.4987562

OpenStudy (amistre64):

crap lol.....n' = 35 ;) e e' = 350

OpenStudy (anonymous):

ok little off track here why are we breaking down the east and north into double derives?

OpenStudy (amistre64):

its hard to keep up with my typos with my computer lagging so much...... you need to read thru them :)

OpenStudy (amistre64):

we derive implicitly with respect to time; so all our derived bits stay intact

OpenStudy (amistre64):

e^2 + n^2 = r^2 ; implicit with respect to "t"

OpenStudy (amistre64):

e' 2e + n' 2n ----------- = r' = dr/dt = how fast the boats r are moving away from each other

OpenStudy (amistre64):

thats 2r under there :) but all the twos factor out to 1 so they can dissapear

OpenStudy (amistre64):

e' e + n' n ----------- = r' = dr/dt r

OpenStudy (anonymous):

fantastic answer as always amistre, thanks a ton for the help

OpenStudy (amistre64):

35(10) + 25(100) ---------------- = r' = dr/dt sqrt(10100)

OpenStudy (amistre64):

youre welcome :)

OpenStudy (amistre64):

2850 ---------- = 28.3585 sqrt(10100)

OpenStudy (anonymous):

yup i agree our math was jiving on that last answer